I need help on this question i tried 2,3,and 4 but are too high."The perimeter of rectangle is 30 inches. If its length is 3 times it's width find the dimensions."
Answer by ewatrrr(24785)

 
Hi,

Let w represent the width
P = 2L + 2w = 30in
0r L + w = 15in
3w + w = 15
4w = 15
w = 15/4 in and length is 45/4 in
Dimensions are
3 3/4 by 11 1/4 inches